X cm³ of 0.15 mol dm³ zinc nitrate solution Zn(NO3)2 is pipetted into a 500 cm³ volumetric flask to produce 500 cm³ of 0.018 mol dm³ zinc nitrate solution. Zn(NO3)2. Determine the value of X.


1
Expert's answer
2021-08-31T01:31:27-0400

Molar mass of Zinc nitrate = 189.36

189.36 × 0.018 =3.41×10 = 34.1

= 500× 0.15/34.1

= 2.2cm3